A circuit consists of 3 resisters of 12,18,and 36 ohms respectvely by join in parallel and the combination is connected in series with a resister of 12 ohms the hole circuit is connected to a 60 v supply calculate current in each branch,total current drown and power dispates in each resister?

Grade:12th pass

1 Answers

Arun
6 years ago
P1 = V^2 / R1, 
P1 = V^2 / 12 = 36 W. 
V^2 = 12*36 = 432, 
V=20.8 Volts. = Voltage across R1,R2,R3 

I1 = V1 / R1 = 20.8 / 12 = 1.73A. 
I2 = 20.8 / 18 = 1.16A. 
I3 = 20.8 / 36 = 0.578A. 

I4=I1 + I2 + I3 = 1.73 + 1.16 + 0.578 = 
3.47A. 

R4=V4 / I4 = (60 - 20.8) / 3.47 = 11.3 
Ohms. 

Pt=E*It = E*I4 = 60 * 3.47 = 208 Watts